Operations Report – February 9th

Crew 221 Operations Report 09Feb2020

SOL: 7

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: on (at 6pm)

Hours run: 11:00

From what time last night: 8:00 pm

To what time this morning: 6:00 am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before the generator is run at night) 80%

Diesel Reading – 97%

Station Propane Reading – 58%

Ethanol Free Gasoline – gallons

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 250 gals

Static to Loft Pump used – yes

Water in GreenHab – 211.6 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: no

Hours: (before EVA) : 129.9

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging)

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: no

Hours: 136.1

Beginning charge: 100%

Ending charge:

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: Suit 10 battery replacement performed.

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment continues, now watered with 4l/day. watering of the plants carried out as per the procedure

Summary of ScienceDome operations: 1 incubator running, microbiology experiments going on and cultures are being grown in agar plates.

Summary of RAM operations: robot assembly is ongoing and expected to continue, antenna assembly for radio communication experiment continued today.

Summary of any observatory issues: nominal

Summary of health and safety issues: Two crew members have a sore throat. Another crew member hit their right knee with a rock during sol 6 EVA

Questions, concerns, and requests to Mission Support: the oven was reconfigured to work with propane as the previous configuration was for natural gas, and that was determined to be the cause of the smell reported. Everything is operational now.

Science Report – February 9th

Crew 221 Science Report 09-FEB-20
Crew Science Officer: Hemani Kalucha

1. Microbiology Research
Today, we put more samples in the incubator to test for bacteria growth. This time, we used our own agar in the samples.

2. ISRU project
Nothing further to report.

3. Mung Beans

Nothing further to report.

4. Radio Communication

Crew Engineer got the antenna to successfully capture and decipher digital signals from around the world (Canada, France, Bangalore!) Photo report shows the transcript of signals captured.

Operations Report – February 9th

Crew 221 Operations Report 07Feb2020

SOL: 7

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: on (at 6pm)

Hours run: 11:00

From what time last night: 8:00pm

To what time this morning: 6:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 80%

Diesel Reading – 97%

Station Propane Reading – 58%

Ethanol Free Gasoline – gallons

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 250 gals

Static to Loft Pump used – yes

Water in GreenHab – 211.6 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: no

Hours: (before EVA) : 129.9

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging)

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: no

Hours: 136.1

Beginning charge: 100%

Ending charge:

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: Suit 10 battery replacement performed.

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment continues, now watered with 4l/day. watering of the plants carried out as per procedure

Summary of ScienceDome operations: 1 incubator runing, microbiology experiments going on and cultures are bieng grown in agar plates.

Summary of RAM operations: robot assembly is ongoing adn expected to continue, antenna asembly for radio comunication experiment continued today.

Summary of any observatory issues: nominal

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support:

Journalist Report – February 9th

Crew 221 Journalist Report 09Feb2020

Author: Hemani Kalucha, Crew Journalist

Sol 4

Title: A little bit of rest

Yesterday evening, we had another amazing cultural night. We learned so much about Costa Rica and found it is such an impressive country. It voluntarily gave up its army in 1948, it has a literacy rate of 97%, and it currently runs on 98% renewable energy. We’ve now decided to have a reunion in Costa Rica next year, especially because of the incredibly yummy food Crew Engineer Luis prepared last night. We also got the chance to try a rare Costa Rican delicacy, the core of a palm tree! This morning, Crew Journalist Hemani made cheesy scrambled eggs for the crew. The crew was served a second luxurious round of coffee and also finally ate the cookies Crew 219 gifted to us as a treat. It was a morning of deliciousness. After an afternoon of great conversation, the crew got busy with planning our EVA experiments and preparing new microbiology samples. In particular, Crew GHO Jessica is trying to plan EVAs for rover operations testing and emergency response testing. She is also trying to use previously published parameters to test EVA success on our EVA operations to compare the results. Tonight, we are all very excited to do the combined Zumba and Salsa sports session prepared by the Commander and Engineer.

Green Hab Report – February 9th

Crew 221 GreenHab Report 09FEB2020

GreenHab Officer: Jessica Todd

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on. Door opened today from 1:05 pm to 4:25 pm.

Average temperatures: 18.33°C

Hours of supplemental light: 19:30 to 00:00

Daily water usage for crops: 14.7 gallons

Daily water usage for research and/or other purposes: 1 gallon

Water in Blue Tank: 211.6 gallons

Time(s) of watering for crops: 2

Changes to crops: Pumpkin plants are flowering nicely. There are a number of soil-filled pots on the bottom level of the central shelf. They are labeled as bell peppers but in one week of monitoring I’ve seen no sprouting. I don’t think there are any plants in these pots.

Narrative: Watered the plants this morning at 9:43am. Today was a little colder so we kept the GreenHab toasty this morning and opened the door for a few hours around lunch. We had omelettes for breakfast so harvested some lovely spinach and rocket to get some greenery! Our mungbeans and blue corn have yet to sprout but will continue to monitor.

Harvest: 9g Chard Spinach, 3g Spinach Lakeside, 6g Rocket

Support/supplies needed: NA

Other Notes: NA

Sol Summary – February 9th

Crew 221 Sol Summary Report 09Feb-2020

Sol:7

Summary Title:

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Update since last Sol summary: The crew had a lovely cultural night of Costa Rica with crew engineer L. Monge. The food he prepared was delicious and he even brought everyone a traditional hat and scarf from his country. The team learned a lot and is planning on a team meeting in Costa Rica once we return to Earth one day.

Today the crew is half way through rotation and therefore took the liberty to sleep one hour more and started its day with delicious scrambled eggs prepared by crew journalist H. Kalucha.

After breakfast, the team sat together, reviewing their performance of the first week. We discussed what we enjoyed and what the team can improve on, which resulted in a 2 hours cleaning activity before each member focused on their science projects.

The microbiology team made great progress, preparing 5 new samples and taking samples from stones directly to see if more bacteria is growing. Most of the samples the team prepared showed growth of bacteria.

Crew engineer L. Monge made good progress on his antenna project, which shall be installed in the field the coming weeks. The crew astronomer G. Alotaibi is still not satisfied with his images and will keep working on it.

As part of the team feels unwell, crew 221 was not on their best performance today and hopes that it will make better progress over the next few days and also should start to do videos for outreach.

Look Ahead Plan: Sports Session, Meditation Session, Planning of Operation EVAs, last microbiology sample collection

Anomalies in work: None

Update mouse in hab: still no joy in catching it.

Weather: cloudy during the day, partially cloudy during the evening light winds

Crew Physical Status: two crew members showing signs of the flu, one crew member injured knee

Reports to be filed:

Sol Summary

Photo Report

Operations Report

GreenHab Report

Science Report

Astronomy Report

EVA Request

HSO Report

Support Requested: Olive Oil, distilled white vinegar

Astronomy Report – February 9th

Astronomy Report

Name: Ghanim Alotaibi

Crew: 221
Date: 09 Feb 2020

Two images were taken last night. One image for V CAS and the other for V0960 MON. Out of the three images submitted for each search pattern, only two were taken.

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested MDRS-14

Objects to be Imaged this Evening:

V CAS (Resubmitted today)
Search pattern (submitted last night): Still active since last night. 1 image is missing out of the three images requested for each search pattern.
V0960 Mon (resubmitted today).

Images submitted with this report: Non

Problems Encountered: No photometry measurement was performed today due to the bad seeing profile (please see an example of the seeing profile for V CAS in the attached file). This seeing profile also occurs with V0960 MON as well as with random stars. The reason of this might be bad focusing and high background noise levels. New requests were submitted, and problem reported to observatory director.

 

Sol Summary – February 9th

Crew 221 Sol Summary Report 09Feb-2020

Sol:7

Summary Title:

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Update since last Sol summary: The crew had a lovely cultural night of Costa Rica with crew engineer L. Monge. The food he prepared was delicious and he even brought everyone a traditional hat and scarf from his country. The team learned a lot and is planning on a team meeting in Costa Rica once we return to Earth one day.

Today the crew is half way through rotation and therefore took the liberty to sleep one hour more and started its day with delicious scrambled eggs prepared by crew journalist H. Kalucha.

After breakfast, the team sat together, reviewing their performance of the first week. We discussed what we enjoyed and what the team can improve on, which resulted in a 2 hours cleaning activity before each member focused on their science projects.

The microbiology team made great progress, preparing 5 new samples and taking samples from stones directly to see if more bacteria is growing. Most of the samples the team prepared showed growth of bacteria.

Crew engineer L. Monge made good progress on his antenna project, which shall be installed in the field the coming weeks. The crew astronomer G. Alotaibi is still not satisfied with his images and will keep working on it.

As part of the team feels unwell, crew 221 was not on their best performance today and hopes that it will make better progress over the next few days and also should start to do videos for outreach.

Look Ahead Plan: Sports Session, Meditation Session, Planning of Operation EVAs, last microbiology sample collection

Anomalies in work: None

Update mouse in hab: still no joy in catching it.

Weather: cloudy during the day, partially cloudy during the evening light winds

Crew Physical Status: two crew members showing signs of the flu, one crew member injured knee

Reports to be filed:

Sol Summary

Photo Report

Operations Report

GreenHab Report

Science Report

Astronomy Report

EVA Request

Support Requested: Olive Oil, distilled white vinegar